OpenStudy (anonymous):

A car with a mass of 103 kilograms traveling at a speed of 50 kilometers/hour (13.9 meters/second) collides with a stationary car. The duration of the collision is 1 second. After the collision the speed of the moving car is reduced to 5 kilometers/hour (1.4 meters/second). What is the value of the force involved in the collision?

OpenStudy (anonymous):

I am looking for an amount, in Newtons. An answer would be great but please explain it as well, or give me a formula?

OpenStudy (shamim):

F*t=mv-mu

OpenStudy (anonymous):

thank you, can you possibly tell me what each of those values represent?

OpenStudy (shamim):

F=force m=103kg u=13.9m/s v=1.4m/s

OpenStudy (shamim):

do u need any more explanation

OpenStudy (anonymous):

it was multiple choice, 1.2 × 10^4 newtons 1.4 × 10^4 newtons 1.6 × 10^3 newtons 1.6 × 10^2 newtons Based on my calculations, I chose 1.4 x 10^4, is that correct?

OpenStudy (shamim):

its midnight here and i lying on my bed. so its not possible to calculate. sorry

OpenStudy (anonymous):

It is okay I got the right answer! thank you so much!

OpenStudy (shamim):

welcome